a swimming pool is 20 feet wide and 40 feet long it is surrounded by square tiles each 1 foot by 1 foot. how many tiles are there surrounding the pool?
Answer by rajagopalan(174)

swimming pool 20 feet wide
and 40 feet long
So Circumference=2(wide+long)
=2(20+40)
=2(60)
=120
So minimum 120 tiles of 1 ft length needed
Plus 4 tiles at the corner
So Totally 120+4=124 tiles
okay
Verify: Take a graph sheet
scale 1ft=1mm
Draw the outline and understand
Answer = 124 tiles